. In natural numbers, x – 5 = – 5 has                                 solution
In natural numbers, x – 5 = – 5 has                                 solution

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the given integers, x–5 =–5.
Transpose - 5 to RHS then it becomes 5.

x = - 5 +5

x= 0
In natural numbers, x–5 =–5 has no solution.

Thanks